OpenStudy (anonymous):

what are the x- and y- intercepts for the graph of 3x+y=15

OpenStudy (noelgreco):

the y-intercept occurs at x = 0, and the x-intercept when y=0.

OpenStudy (anonymous):

?

OpenStudy (gorv):

ato find x intercept put y=0 and calculate x

OpenStudy (gorv):

that is 3x+0=15

OpenStudy (gorv):

now solve it further

OpenStudy (anonymous):

5?

OpenStudy (gorv):

yeah that is x intercept

OpenStudy (anonymous):

k so now what about the y

OpenStudy (gorv):

now for y intercept ..at y intercept x=0

OpenStudy (gorv):

so put x=0 for y intercept

OpenStudy (gorv):

3*0+y=15 0+y=15

OpenStudy (gorv):

what u got for y??

OpenStudy (anonymous):

x is 3 and y is 15

OpenStudy (gorv):

yep they are your intercept